# UDC-VIT: A Real-World Video Dataset for Under-Display Cameras

### v1

@@ Appendix E Reproducibility
The learnable restoration models used for evaluating the UDC-VIT dataset include DISCNet (Feng et al., 2021), UDC-UNet (Liu et al., 2022b), FastDVDNet (Tassano et al., 2020), EDVR (Wang et al., 2019), ESTRNN (Zhong et al., 2020), and DDRNet (Liu et al., 2024). We use a single-node GPU cluster to train each benchmark model. Each node has eight AMD Instinct MI100 GPUs. While we mainly stick to the original authors’ code and training settings for the models, we introduce some modifications except ESTRNN.
  • DISCNet. DISCNet is designed to restore UDC still images in high dynamic range (HDR) (e.g., SYNTH (Feng et al., 2021)). Accordingly, we modify the PyTorch DataLoader to use normalization instead of Reinhard tone mapping (Reinhard et al., 2002). The DataLoader randomly selects one frame per video from the UDC-VIT dataset for each iteration during the training and validation phases.
  • UDC-UNet. UDC-UNet is also designed to restore UDC still images in HDR. The original authors do not conduct normalization or tone mapping in the DataLoader and employ a tone mapping L1 loss function. However, since the UDC-VIT dataset has a low dynamic range (LDR), we modify the PyTorch DataLoader to use normalization. We clamp the model output between 0 and 1 and then calculate the L1 loss. The DataLoader randomly selects one frame per video from the UDC-VIT dataset for each iteration.
  • FastDVDNet. FastDVDNet is a video denoising model that utilizes NVIDIA’s Data Loading Library (DALI) (Nvidia, 2018), processing a noise map and multiple frames as inputs. Instead of DALI, we employ the PyTorch DataLoader tailored to the UDC-VIT dataset in npy format. We set the noise level to zero. To accommodate FHD resolution and multiple degradations in the UDC-VIT dataset, we increase the patch size from 64 to 256. Furthermore, we extend the training duration of FastDVDNet to 400 epochs, compared to the original 95, to ensure the model reaches full saturation.
  • EDVR. To address out-of-memory issues with EDVR, which boasts 23.6 M parameters, we reduce the patch size from 256 to 192. Additionally, during inference on the test set, we divide it into two patches of size $3\times 1,060\times 1,060$ each and merge them afterward.
  • DDRNet. During the inference process, the authors of DDRNet partition each frame into patches of size $3\times 256\times 256$ and input 50 frames simultaneously. However, patch-wise inference introduces the borderline between patches. To address this, we conduct inference at full resolution ($3\times 1,060\times 1,900$) with ten frames at a time.
@@ Appendix F Discussion on the Responsible Use of the Dataset
This section discusses the potential negative societal impacts, the corresponding user guidelines, and our responsibility.
@@ F.1 Potential Negative Societal Impacts
The UDC-VIT dataset includes the faces and motions of 22 research participants, raising concerns about its potential for misuse, such as in deep fake applications. This technology can generate convincingly altered videos, threatening individual privacy and societal trust. Deep fakes can infringe upon personal integrity and privacy, leading to social unrest and confusion. Given these potential negative societal impacts, careful consideration is needed when using the dataset.
@@ F.2 User Guidelines
The users of the UDC-VIT dataset are expected to adhere to the following guidelines:
  • Responsible use. Users must ethically and responsibly utilize the dataset, ensuring it does not infringe on individual privacy or contribute to societal harm.
  • Compliance with legal and ethical standards. Users must comply with all relevant legal and ethical standards. They should obtain Institutional Review Board (IRB) approvals by the regulations of their respective countries and respect any restrictions or conditions imposed by the IRB or other regulatory bodies. Any violation of the laws will be the user’s sole responsibility.
  • Restricted Usage. Users must avoid using the UDC-VIT dataset for harmful applications, such as deep fake technologies or other misinformation or manipulation. Furthermore, the 22 participants’ agreed-upon research scope during our IRB review centers on acquiring UDC video datasets and developing restoration models. Therefore, this dataset must exclusively serve UDC research purposes.
@@ F.3 Our Responsibility
  • Protect participant privacy. Our primary concern is to preserve the privacy and confidentiality of research participants. Although participants consented to the public use of their faces and motions within the dataset, we are dedicated to providing user guidance for appropriate research utilization and exerting efforts to safeguard other personal information.
  • Facilitate ethical use. We provide comprehensive guidelines and documentation on datasheets for datasets, our project site, and our research group’s website. The email automatically sends the download link when users complete the application form on our research group’s website, which will also provide information about the dataset’s potential risks and ethical considerations.
  • Respond to concerns. Our commitment to responsible management of the UDC-VIT dataset extends to responding promptly to any concerns or complaints raised. We value user feedback and are ready to take appropriate actions, such as data corrections and updates, to mitigate potential harm or misuse if any misuse of the dataset is reported on the website of our research group, as shown in Figure A.3.

### v2

@@ Appendix E Reproducibility
The learnable restoration models used for evaluating the UDC-VIT dataset include DISCNet [18], UDC-UNet [33], FastDVDNet [54], EDVR [57], ESTRNN [62], and DDRNet [31]. We use a single-node GPU cluster to train each benchmark model. Each node has eight AMD Instinct MI100 GPUs. While we mainly stick to the original authors’ code and training settings for the models, we introduce some modifications to all models except ESTRNN.
  • DISCNet. DISCNet is designed to restore UDC still images in high dynamic range (HDR) (e.g., SYNTH [18]). Accordingly, we modify the PyTorch DataLoader to use normalization instead of Reinhard tone mapping [40]. The DataLoader randomly selects one frame per video from the UDC-VIT dataset for each iteration during the training and validation phases.
  • UDC-UNet. UDC-UNet is designed to restore UDC still images in HDR. The original authors do not conduct normalization or tone mapping in the DataLoader and employ a tone mapping L1 loss function. However, since the UDC-VIT dataset has a low dynamic range (LDR), we modify the DataLoader to normalize the input. The model output is clamped between 0 and 1 before computing the L1 loss. The DataLoader randomly selects one frame per video from the UDC-VIT dataset for each iteration.
  • FastDVDNet. FastDVDNet is a video denoising model that utilizes NVIDIA’s Data Loading Library (DALI) [36], processing a noise map and multiple frames as inputs. Instead of DALI, we employ the PyTorch DataLoader tailored to the UDC-VIT dataset in NPY format. We set the noise level to zero. To accommodate the FHD resolution and multiple degradations in the UDC-VIT dataset, we increase the patch size from 64 to 256. Moreover, we extend the training duration of FastDVDNet from the original 95 epochs to 400, allowing the model to reach full convergence.
  • EDVR. To address the out-of-memory issues of EDVR, which contains 23.6M parameters, we reduce the training patch size from 256 to 192. During inference on the test set, we divide each frame into two patches of size 3×1060×10603$\times$ 1060$\times$ 10603 × 1060 × 1060 and merge the outputs afterward.
  • DDRNet. During inference, the authors of DDRNet partition each frame into patches of size 3×256×2563$\times$ 256$\times$ 2563 × 256 × 256 and process 50 frames simultaneously. However, this patch-wise inference introduces visible seams at the patch boundaries. To mitigate this artifact, we instead perform inference at full resolution (3×1060×19003$\times$ 1060$\times$ 19003 × 1060 × 1900), using ten consecutive frames at a time.
@@ Appendix F Discussion on the Responsible Use of the Dataset
This section discusses potential negative societal impacts, corresponding user guidelines, and our responsibilities.
@@ F.1 Potential Negative Societal Impacts
The UDC-VIT dataset contains facial images and motions of 22 research participants, which raises concerns about potential misuse, such as in deepfake applications. This technology can generate convincingly altered videos, posing risks to individual privacy and societal trust. Deepfakes may infringe upon personal integrity and privacy, potentially leading to social unrest and confusion. Given these potential negative societal impacts, careful consideration is required when using the dataset.
@@ F.2 User Guidelines
Users of the UDC-VIT dataset are expected to adhere to the following guidelines:
  • Responsible use. Users must utilize the dataset ethically and responsibly, ensuring that its use does not cause societal harm or infringe on individual privacy.
  • Compliance with legal and ethical standards. Users must comply with all applicable legal and ethical standards. This includes obtaining approval from an Institutional Review Board (IRB) by the regulations of the respective country and adhering to any restrictions or conditions imposed by the IRB or other regulatory bodies. Any legal violations will be the sole responsibility of the user.
  • Restricted usage. The UDC-VIT dataset must not be used for harmful applications, such as deep fake technologies or the creation of misinformation and manipulation. Furthermore, the 22 participants agreed to a research scope defined during our IRB review, which is limited to acquiring UDC video data and developing restoration models. Therefore, the dataset must be used exclusively for UDC-related research purposes.
@@ F.3 Our Responsibility
  • Protect participant privacy. Our primary priority is to maintain the privacy and confidentiality of our research participants. Although the participants provided consent for the public use of their facial appearances and movements within the dataset, we are committed to guiding users toward appropriate research practices and to making efforts to safeguard any additional personal information.
  • Facilitate ethical use. We provide comprehensive documentation and ethical usage guidelines through the Datasheets for Datasets and our GitHub repository at https://github.com/mcrl/UDC-VIT. The dataset can be accessed via an automatic email system after users complete the application form on our research group’s website. This process also informs users of potential risks and ethical considerations associated with the dataset.
  • Respond to concerns. We are committed to the responsible stewardship of the UDC-VIT dataset and will respond promptly to any concerns or complaints that may arise. We value user feedback and are prepared to take appropriate actions—such as data correction or updates—to prevent or mitigate harm, should any misuse of the dataset be reported through our research group’s website (see Fig. D.3).